OpenStudy (anonymous):

Airplane A travels 1400 km at a certain speed. Plane B travels 1000 km at a speed 50 km/h faster than plane A in 3 hrs less time. Find the speed of each plane.

OpenStudy (anonymous):

let da = 1400 km let Va = speed of plane A let ta = time a let Vb = speed of plane B let db = 1000 km let tb = time b db = Vb*tb 1000 km = (Va + 50 km/h)(da/Va - 3h) 1000 km = (Va + 50 km/h)(1400km/Va - 3h) now solve for Va add 50 km/h to get Vb

OpenStudy (anonymous):

oh a relationship between da, Va and ta may be helpful da = Va*ta db = Vb*tb

OpenStudy (anonymous):

did this help?